OpenStudy (anonymous):

Carbon-14 has a half-life of 5,730 years. In how many years will 120 grams of Carbon-14 decay to 15 grams? A. 11,460 years B .17,190 years C. 22, 920 years D. 28,650 years

OpenStudy (anonymous):

the radioactive decay equation: \[N(t)=N_0e^{-t/\tau}\\ \implies {N\over N_0}=e^{-t/\tau}\\ \ln\left(N\over N_0\right)={-t\over \tau}\\ t = \tau \ln\left(N_0\over N\right)\\ \;\\ N=15\,{\rm g}\\ N_0=120\,{\rm g}\\ \tau=5730\\ t=? \]